Session: May 2000 (session 2)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Huang Xinyong at Tue May 30 13:55:31 METDST 2000) Success
The tape recorder has trouble from UT: 11:34:04 to 11:36:49. The error is: "the rc derection of tape travel is different than comanded."

Medicina (Maccaferri A. at Tue May 30 14:09:09 METDST 2000) Success
Everything ok. The weather was raining.

Hartebeesthoek (Jonathan Quick at Tue May 30 14:21:35 METDST 2000) Success
No problems

Effelsberg (A. Kraus at Tue May 30 14:25:57 METDST 2000) Success

Onsala (A. Polatidis at Tue May 30 15:53:04 METDST 2000) Success

Ny_Alesund (David Holland at Tue May 30 16:23:21 METDST 2000) Minor failures
We found a problem with the upper tape reel causing the tape to moove slightly when coming off the reel. This was rectified after stopping the tape for a short time at 09.32.

Wettzell (e.bauernfeind at Tue May 30 17:49:34 METDST 2000) Success
wettzell observed c002xwz successful. frequency standard: maser efos-13 delta ut(c;wz to efos-13)=+6.57 microsec. (151/08:45 ut) delta ut(c;wz to efos-13)=+6.60 microsec. (151/14:45 ut) delta ut(c;wz to tac2) =+2.95 microsec. (151/08:45 ut) delta ut(c;wz to tac2) =+2.93 microsec. (151/14:45 ut) wx/covered

Urumqi (Liu X. at Wed May 31 06:45:57 METDST 2000) Minor failures
The head calibration problems.

Yebes (Paco Colomer at Thu Jun 1 14:37:32 METDST 2000) Success

Noto (C.Trigilio at Fri Jun 2 12:28:25 METDST 2000) Success
No one apparent problem

Urumqi (Liu X. at Mon Jun 5 07:58:46 METDST 2000) Minor failures
The auto head caliration can not be done, the absolute offset of read is set to 600 microns in head.ctl. The read is very bad, but data maybe writen on the tape. The Tsys in vc02 and vc04 is not correct, is should be 110K.

Westerbork (Tony Foley at Mon Jun 5 17:15:29 METDST 2000) Minor failures
3.6cm only. Half the BBCs with LCP and half RCP. Used 12 telescopes in the tied array